OpenStudy (mathhelp346):

A boat went 75 miles in 2 hours with the current and 4 hours without. What was the boat and current's speed?

OpenStudy (anonymous):

d=vt v=d/t v(boat)=75mi/4hr=18.75 mi/hr v(boat)+v(current)=75mi/2hr 18.75mi/hr +v(current)=75mi/2hr v(current)=75mi/2hr-18.75mi/hr=18.75mi/hr

OpenStudy (mathhelp346):

whats v

OpenStudy (anonymous):

velocity

OpenStudy (mathhelp346):

ok

OpenStudy (mathhelp346):

so the speed of the boat is 18.75?

OpenStudy (anonymous):

yes the boat and the current travel at the same speed, which is why the time doubles when the current is not flowing

OpenStudy (mertsj):

Are you sure you posted the problem correctly? Typically these problems have the boat going with the current and against the current.

OpenStudy (mathhelp346):

yea thats what is says

OpenStudy (mertsj):

Ok. Good. Thanks for checking.

OpenStudy (mathhelp346):

i keep getting 28.125 for the boats speed

OpenStudy (mertsj):

The boat went 75 miles in 4 hours without any current. So its speed must be 75/4=18.75

OpenStudy (mathhelp346):

oh ok